# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 5. RQ3: What techniques are used to optimize and evaluate LLM4SE?
5.1. What optimizers are used to enhance model performance?
We examined the parameter and learning rate optimizers reported in our selected primary studies. As shown in Fig. 7 (a), fine-tuning emerges as the most widely used optimization algorithm in LLM studies, appearing in 87 research works (Niu et al., 2022; Thapa et al., 2022; Von der Mosel et al., 2022; Wang et al., 2023d; Zheng et al., 2023b; Al-Kaswan et al., 2023; Deng et al., 2023d; Döderlein et al., 2022; Jiang et al., 2023b). This signifies the dominance of fine-tuning in adapting pre-trained models to specific tasks, resulting in enhanced performance across various natural language processing tasks (Cassano et al., 2023; Chen et al., 2023c; Ciniselli et al., 2021). Hyperparameter optimization is another prominent approach, found in 55 studies (Ochs et al., 2023; Patil et al., 2023; Saieva et al., 2023; Tufano et al., 2022; Wan et al., 2022b), highlighting its significance in fine-tuning the hyperparameters of language models to achieve optimal performance on specific tasks (Zhang et al., 2023c; Zan et al., 2022b; Yang et al., 2022c). Of the data described above, 35 of these studies (Zhang et al., 2023c; Zhu et al., 2023) used both fine-tuning and hyperparameter parameter optimization algorithms. The limited occurrences of Bayesian (Prenner and Robbes, 2021) and Stochastic Gradient Descent (SGD) (Nasir et al., 2023) optimization (one study each) suggest that they are less frequently employed in LLM research.
Among the learning rate optimization algorithms illustrated in Fig. 7 (b), Adam stands out with 25 occurrences in the studies (Chen et al., 2023a; Ciborowska and Damevski, 2023; He et al., 2023; Huang et al., 2023b; Kolthoff et al., 2023; Kou et al., 2023a). Adam is an adaptive optimization algorithm that combines adaptive learning rates with momentum, facilitating faster convergence and reducing the risk of getting stuck in local minima during training (Mirzadeh et al., 2020). Similarly, AdamW appears in 21 studies (Deng et al., 2023c; Fatima et al., 2022; Fu and Tantithamthavorn, 2022; Hey et al., 2020; Khan et al., 2021), demonstrating its significance in improving generalization by adding weight decay regularization to Adam (Yuan et al., 2022). ZeRO (Hendrycks et al., 2021; Thapa et al., 2022) and Adafactor (Gupta et al., 2023; Ye et al., 2023) are relatively less explored, mentioned in three and two studies respectively. NVLAMB is found in one study (Von der Mosel et al., 2022), indicating limited exploration of this specific learning rate optimization algorithm in LLM research.
Figure 7. Various optimizers used in the collected studies.
@@ 5.2. What prompt engineering techniques are applied to improve the performance of LLMs in SE tasks?
Large-scale pre-trained models have demonstrated effectiveness across numerous code intelligence tasks. These models are initially pre-trained on extensive unlabeled corpora and subsequently fine-tuned on downstream tasks. However, the disparity in input formats between pre-training (Ahmad et al., 2021) and downstream tasks (Devlin et al., 2018; Liu et al., 2016) poses challenges in fully harnessing the knowledge embedded in pre-trained models. Furthermore, the efficacy of fine-tuning (Kanade et al., 2020) strongly hinges on the volume of downstream data (Gu et al., 2021; Han et al., 2022; Lester et al., 2021), a circumstance frequently characterized by data scarcity (Feng et al., 2020; Guo et al., 2020; Wang et al., 2021a).
Recent research in the domain of Natural Language Processing (NLP) underscores that prompt engineering (Feng and Chen, 2023; Liu et al., 2023f), as an emerging fine-tuning (Lester et al., 2021; Li and Liang, 2021) paradigm, holds the potential to mitigate the aforementioned issues, yielding commendable outcomes across diverse NLP tasks. In the study conducted by Wang et al. (Wang et al., 2022d), they delve into the application of prompt engineering techniques to enhance the performance of LLMs in SE tasks. The research chiefly explores two prompt types: hard prompts (Gu et al., 2021; Han et al., 2022; Iyer et al., 2016) and soft prompts (Li and Liang, 2021; Han et al., 2022; Tsimpoukelli et al., 2021). Hard prompts entail manually predefined natural language instructions, while soft prompts, including plain soft prompts, replace natural language tokens in hard prompts with surrogate tokens. A variant known as prefix soft prompts adds a few surrogate tokens before the original input. In the context of prompt engineering, the task-specific knowledge imparted by inserted prompts is particularly advantageous for tasks characterized by data scarcity.
Prompt engineering, characterized by the careful design of specialized prompts, has become a fundamental technique for enhancing interactions with LLMs such as ChatGPT (Dong et al., 2023; White et al., 2023b) and WizardCoder (Luo et al., 2023). These tailored prompts serve dual purposes: they direct the LLMs toward generating specific outputs and also function as an interface to access the extensive knowledge embedded in these models. This becomes particularly important in scenarios where traditional datasets, such as those derived from software repositories or standard benchmarks, are either limited or lack the granularity required for specific tasks. In the context of LLM4SE — which encompasses a range of tasks such as code generation (Bareiß et al., 2022; Li et al., 2023c; Tan et al., 2023; Tian et al., 2020), code summarization (Gao et al., 2023; Ahmed et al., 2023), program repair (Huang et al., 2023c; Paul et al., 2023b; Xia and Zhang, 2023a; Yuan et al., 2022), and test generation (Siddiq et al., 2023b; Vikram et al., 2023; Xie et al., 2023) — the role of prompt engineering is indispensable. In summary, recent research highlights the critical role of prompt engineering in enhancing the performance of LLMs for targeted SE tasks, thereby contributing to the evolution of automated software development methodologies.
@@ 5.3. How are evaluation metrics utilized to assess the performance of LLM4SE tasks?
Problem Type Metric # Studies References
Regression MAE (Mean Absolute Error) 1 (Fu and Tantithamthavorn, 2022)
Classification F1-score 21 (Alhamed and Storer, 2022) (Biswas et al., 2020) (Chen et al., 2023a) (Ezzini et al., 2022) (Fatima et al., 2022) (He et al., 2022) (Hey et al., 2020) (Huang et al., 2023b) (Khan et al., 2021) (Khan and Uddin, 2022) (Koide et al., 2023) (Kolthoff et al., 2023) (Le-Cong et al., 2023) (Luo et al., 2022) (Mukherjee and Hellendoorn, 2023) (Sghaier and Sahraoui, 2023) (Sharma et al., 2022) (Thapa et al., 2022) (Zhou et al., 2023b) (Yang et al., 2022c) (Zhang et al., 2020b)
Precision 20 (Biswas et al., 2020) (Chen et al., 2023a) (Dou et al., 2023) (Ezzini et al., 2022) (Fatima et al., 2022) (He et al., 2022) (Hey et al., 2020) (Huang et al., 2023b) (Khan et al., 2021) (Khan and Uddin, 2022) (Koide et al., 2023) (Kolthoff et al., 2023) (Le-Cong et al., 2023) (Mukherjee and Hellendoorn, 2023) (Sghaier and Sahraoui, 2023) (Sharma et al., 2022) (Thapa et al., 2022) (Tian et al., 2023a) (Yang et al., 2022c) (Zhang et al., 2020b)
Recall 18 (Biswas et al., 2020) (Chen et al., 2023a) (Dou et al., 2023) (Ezzini et al., 2022) (Fatima et al., 2022) (He et al., 2022) (Hey et al., 2020) (Huang et al., 2023b) (Khan et al., 2021) (Khan and Uddin, 2022) (Koide et al., 2023) (Kolthoff et al., 2023) (Le-Cong et al., 2023) (Mukherjee and Hellendoorn, 2023) (Sghaier and Sahraoui, 2023) (Sharma et al., 2022) (Tian et al., 2023a) (Zhang et al., 2020b)
Accuracy 12 (Chen et al., 2023a) (Gomes et al., 2023) (Jesse et al., 2022) (Kanade et al., 2020) (Khan et al., 2021) (Khan and Uddin, 2022) (Koide et al., 2023) (Le-Cong et al., 2023) (Lee et al., 2022) (Manh et al., 2023) (Saberi et al., 2023b) (Tian et al., 2023a)
ROC (Receiver Operating Characteristic) 1 (Alhamed and Storer, 2022)
AUC (Area Under the ROC Curve) 4 (Alhamed and Storer, 2022) (Tian et al., 2023a) (Wang et al., 2023d) (Yang et al., 2022c)
FPR (False Positive Rate) 4 (Chen et al., 2023a) (Sun et al., 2023) (Thapa et al., 2022) (Wang et al., 2023d)
FNR (Falser Negative Rate) 3 (Sun et al., 2023) (Thapa et al., 2022) (Wang et al., 2023d)
MCC (Matthews Correlation Coefficient) 1 (Yang et al., 2022c)
Recommendation MRR (Mean Reciprocal Rank) 9 (Ciborowska and Damevski, 2022) (Izadi et al., 2022) (Li et al., 2021) (Lin et al., 2021) (Luitel et al., 2023) (Rahmani et al., 2023) (Salza et al., 2022) (Shi et al., 2022) (Wei et al., 2022a)
Precision@k 4 (Ciborowska and Damevski, 2022) (He et al., 2023) (Lin et al., 2021) (Zhu et al., 2023)
F1-score@k 4 (He et al., 2023) (Lin et al., 2021) (Zhu et al., 2023) (Zhu et al., 2022)
MAP/MAP@k 3 (Ciborowska and Damevski, 2022) (Isotani et al., 2021) (Lin et al., 2021)
Accuracy 3 (Izadi et al., 2022) (Li et al., 2021) (Salza et al., 2022)
Recall@k 2 (He et al., 2023) (Zhu et al., 2023)
Generation BLEU/BLEU-4/BLEU-DC 29 (Ahmed et al., 2023) (Al-Kaswan et al., 2023) (Arakelyan et al., 2023) (Chen et al., 2022) (Ciniselli et al., 2021) (Dibia et al., 2022) (Gao et al., 2023) (Geng et al., 2024) (Jana et al., 2023) (Kuznia et al., 2022) (Li et al., 2023d) (Li et al., 2023a) (Liu et al., 2023a) (Mastropaolo et al., 2021a) (Mastropaolo et al., 2022b) (Mastropaolo et al., 2021b) (Murali et al., 2023) (Niu et al., 2022) (Paul et al., 2023a) (Saberi et al., 2023a) (Shirafuji et al., 2023) (Tufano et al., 2022) (Wang et al., 2023c) (Weyssow et al., 2023) (Yang et al., 2023c) (Ye et al., 2023) (Zan et al., 2023a) (Zhang et al., 2023f) (Zheng et al., 2023b)
Pass@k 28 (Bui et al., 2023) (Cassano et al., 2023) (Chen et al., 2023c) (Chen et al., 2021b) (Dibia et al., 2022) (Döderlein et al., 2022) (Dong et al., 2023) (Du et al., 2023) (Fakhoury et al., 2023) (Jiang et al., 2023b) (Jiang et al., 2023a) (Lahiri et al., 2022) (Li et al., 2023b) (Li et al., 2023e) (Liu et al., 2023c) (Luo et al., 2023) (Olausson et al., 2023) (Shirafuji et al., 2023) (Thakur et al., 2023) (Wang et al., 2023e) (Weyssow et al., 2023) (Yu et al., 2023) (Zan et al., 2022a) (Zan et al., 2022b) (Zan et al., 2023a) (Zhang et al., 2023g) (Zhang et al., 2023e) (Zheng et al., 2023b)
Accuracy/Accuracy@k 16 (Feng and Chen, 2023) (Huang et al., 2023d) (Jain et al., 2022) (Jin et al., 2023) (Kang et al., 2022) (Lahiri et al., 2022) (Liu et al., 2020) (Mastropaolo et al., 2022b) (Mastropaolo et al., 2021b) (Niu et al., 2022) (Pegolotti et al., 2023) (Rao et al., 2023a) (Schlag et al., 2023) (Shirafuji et al., 2023) (Tian et al., 2020) (Zhang et al., 2022a)
CodeBLEU 12 (Arakelyan et al., 2023) (Gao et al., 2023) (Jana et al., 2023) (Jiang et al., 2023a) (Liu et al., 2023a) (Paul et al., 2023a) (Shirafuji et al., 2023) (Wang et al., 2023c) (Weyssow et al., 2023) (Zan et al., 2023a) (Zhang et al., 2023f) (Zheng et al., 2023b)
EM (Exact Match) $1$ (Al-Kaswan et al., 2023) (Gao et al., 2023) (Gupta et al., 2023) (Jana et al., 2023) (Murali et al., 2023) (Paul et al., 2023a) (Shirafuji et al., 2023) (Wang et al., 2023c) (Weyssow et al., 2023) (Ye et al., 2023) (Zhang et al., 2023c)
ROUGE/ROUGE-L 10 (Ahmed et al., 2023) (Al-Kaswan et al., 2023) (Gao et al., 2023) (Geng et al., 2024) (Li et al., 2022f) (Li et al., 2023a) (Mastropaolo et al., 2021b) (Mastropaolo et al., 2022b) (Niu et al., 2022) (Zan et al., 2023a)
METEOR 6 (Ahmed et al., 2023) (Al-Kaswan et al., 2023) (Chen et al., 2022) (Gao et al., 2023) (Geng et al., 2024) (Niu et al., 2022)
Precision 3 (Kou et al., 2023a) (Tian et al., 2020) (Wang et al., 2022c)
Recall 3 (Kou et al., 2023a) (Tian et al., 2020) (Wang et al., 2022c)
F1-score 3 (Kou et al., 2023a) (Tian et al., 2020) (Wang et al., 2022c)
MRR (Mean Reciprocal Rank) 2 (Chen et al., 2022) (Niu et al., 2022)
MFR (Mean First Ranking) 1 (Tu et al., 2023)
MAR (Mean Average Ranking) 1 (Tu et al., 2023)
ES (Edit Similarity) 1 (Liu et al., 2023e)
PP (Perplexity) 1 (Xu et al., 2022)
Evaluating the performance of LLM4SE is a crucial aspect of their development and deployment (Kang et al., 2022). Benchmarking against existing datasets and using baselines are common practices to evaluate the effectiveness of LLMs (Cassano et al., 2023). However, given the diversity of SE tasks, a single evaluation metric may not suffice to capture the model’s performance comprehensively. Thus, researchers often employ a range of evaluation metrics tailored to specific problem types (Mastropaolo et al., 2021b; Niu et al., 2022; Salza et al., 2022). We categorize the SE tasks summarized from 229 papers into four categories according to their addressed problem types, i.e., regression, classification, recommendation, and generation tasks, as displayed in Fig. 8 (b). The selection of evaluation metrics depends on the target problem types. For example, MAE (Mean Absolute Error) has been used for regression tasks (Fu and Tantithamthavorn, 2022). We summarize the most frequently used evaluation metrics for each task type.
@@
For classification tasks, the most commonly used metrics are F1-score (Alhamed and Storer, 2022; Biswas et al., 2020; Chen et al., 2023a; Ezzini et al., 2022; Fatima et al., 2022; He et al., 2022), Precision (Biswas et al., 2020; Chen et al., 2023a; Ezzini et al., 2022; Fatima et al., 2022; He et al., 2022), and Recall (Biswas et al., 2020; Chen et al., 2023a; Ezzini et al., 2022; Fatima et al., 2022; He et al., 2022; Hey et al., 2020), with 21, 20, and 18 studies, respectively, employing these metrics. For example, in the study conducted by Khan et al. (Khan et al., 2021), F1-score is utilized to evaluate the performance of an automatic bug-fixing model. Similarly, Sharma et al. (Sharma et al., 2022) use Precision and Recall to assess the effectiveness of a transformer-based model for code summarization. These metrics are essential for evaluating the model’s ability to correctly classify code snippets (Fatima et al., 2022) or identify specific SE properties (Chen et al., 2023a).
@@
For recommendation tasks, MRR (Mean Reciprocal Rank) is the most frequent metric, used in 9 studies (Ciborowska and Damevski, 2022; Izadi et al., 2022; Li et al., 2021; Lin et al., 2021; Rahmani et al., 2023; Salza et al., 2022; Shi et al., 2022; Wei et al., 2022a). MRR is employed to measure the effectiveness of recommendation systems for code completion, as demonstrated in the study by Ciborowska et al. (Ciborowska and Damevski, 2022). Precision@k (He et al., 2023; Ciborowska and Damevski, 2022; Lin et al., 2021; Zhu et al., 2023) and F1-score@k (He et al., 2023; Lin et al., 2021; Zhu et al., 2022, 2023) are also utilized in recommendation tasks, with 4 studies each. These metrics are used to evaluate the precision and F1-score of the recommended code snippets or code completions.

### v6

@@ 5. RQ3: What techniques are used to optimize and evaluate LLM4SE?
5.1. What tuning techniques are used to enhance the performance of LLMs in SE tasks?
Through surveying research related to LLM4SE, we found that while many general-purpose LLMs (e.g., ChatGPT) can be directly applied to software engineering tasks such as code generation (Dong et al., 2023b; Liu et al., 2023a; Yetiştiren et al., 2023), code summarization (Shi et al., 2023c; Sun et al., 2023b; Yang et al., 2023d), and program repair (Charalambous et al., 2023; Gao et al., 2023b; Xia and Zhang, 2023b) without fine-tuning, the hidden potential of LLMs often needs to be realized through tuning to be fully exploited. Specifically, this requires training LLMs with task-specific data to learn knowledge relevant to the task context to perform better. We observed that out of 83 studies, LLMs were fine-tuned using full fine-tuning techniques to adapt to downstream SE tasks, with the majority being BERT series models (Ciniselli et al., 2021; Ezzini et al., 2022; Fatima et al., 2022; Jesse et al., 2022; Kou et al., 2023a; Lee et al., 2022; Lin et al., 2021; Luo et al., 2022; Salza et al., 2022; Von der Mosel et al., 2022; Wang et al., 2022c; Wei et al., 2022a; Zhang et al., 2022a). The cost of training these LLMs is expensive, requiring a large amount of computational resources and massive amounts of data. It is also costly to train and deploy the fine-tuned models separately for each downstream task, as the traditional fine-tuning approach would need to copy a model and perform full-parameter fine-tuning for each downstream task (Cassano et al., 2023; Deng et al., 2023c; Ezzini et al., 2022; Izadi et al., 2022; Jesse et al., 2022; Lee et al., 2022).
To reduce this computational burden, some researchers have previously used In-Context Learning (ICL) (Gao et al., 2023b; Geng et al., 2024; Hu et al., 2024; Huang et al., 2023f; Jiang et al., 2023a), which feeds the model with manually designed “prompts” that are overly reliant on human design and do not require updating model parameters at all. However, ICL only operates at the time of inference and does not involve learning task-specific parameters, which experimentally proved to give the model limited improvement in downstream tasks (Liu et al., 2022a). To address this problem, researchers have begun to apply Parameter Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019) techniques to LLMs. PEFT aims to improve the performance of pre-trained models on new tasks by optimizing the subset of parameters fine-tuned, thereby reducing the overall computational complexity. This approach maintains the majority of the pre-trained model’s parameters in a fixed state, focusing fine-tuning efforts on a minimal yet impactful set of parameters (Weyssow et al., 2023a). Prior code intelligence research has demonstrated the capabilities of PEFT techniques, frequently revealing their superiority over full fine-tuning on a variety of tasks (Weyssow et al., 2023a). Four common techniques of PEFT include Low-Rank Adaptation (LoRA) (Hu et al., 2021), prompt tuning (Lester et al., 2021), prefix tuning (Li and Liang, 2021), and adapter tuning (Houlsby et al., 2019). We now elaborate on each as follows:
Low-Rank Adaptation (LoRA). LoRA injects low-rank trainable matrices into the attention layers of the Transformer architecture to significantly reduce the number of trainable parameters. We observed that eight studies (Arakelyan et al., 2023; Lu et al., 2023; Pan et al., 2023c; Shestov et al., 2024; Shi et al., 2023c; Silva et al., 2023; Wang et al., 2023e; Zhang et al., 2023k) utilized LoRA to enhance the performance of LLMs in SE tasks. For instance, Pan et al. (Pan et al., 2023c) trained SteloCoder, specifically designed for translating multiple programming languages into Python code, which is based on the StarCoder LLM. LoRA technology was employed during the modification of the StarCoder model architecture to adjust the parameter count. Additionally, Silva et al. (Silva et al., 2023) applied LoRA to LLaMA, resulting in a highly effective “program repair adapter” for fixing bugs through fine-tuning.
Prompt tuning. Prompt tuning involves appending learnable tokens to the model’s input, guiding it towards better task performance. This method keeps the model’s architecture unchanged, leveraging adaptable prompts to influence outputs without altering internal parameters. In the surveyed papers, three research works (Lu et al., 2023; Wang et al., 2023e; Zhu et al., 2023) utilized prompt tuning. For instance, Zhu et al. (Zhu et al., 2023) proposed a method named AUMENA, which automates method naming tasks through context-aware prompt tuning.
Prefix tuning. Prefix tuning adapts pre-trained language models by adding trainable tokens not just to the input but also across internal layers, affecting the model’s intermediate representations. This approach modifies the model’s processing with minimal changes to its original parameters, allowing for task-specific customization. This technique was utilized in the following two studies: Lu et al. (Lu et al., 2023) fine-tuned LLaMA-Reviewer for automating code review, while Wang et al. (Wang et al., 2023e) fine-tuned CodeT5+ for multiple downstream tasks such as code completion, code generation, and code search.
Adapter tuning. Adapter tuning adds small neural network modules to the original model, then fine-tuning them on specific tasks without altering the original model’s parameters. Agarwal et al. (Agarwal et al., 2024) fine-tuned LLMs using adapter tuning techniques to make them suitable for code representation tasks. Wang et al. (Wang et al., 2023a) indicated that LLMs refined through adapter tuning perform exceptionally well in code search and code summarization tasks.
In addition to the above-mentioned tuning methods, other techniques have been used for tuning LLMs in the LLM4SE domain, such as Reinforcement Learning (RL) (Islam and Najafirad, 2024; Islam et al., 2024; Jain et al., 2023a; Steenhoek et al., 2023; Yang et al., 2023f), Supervised Fine Tuning (SFT) (Dong et al., 2023c; Islam and Najafirad, 2024; Mastropaolo et al., 2022b; Steenhoek et al., 2023; Yang et al., 2023f), an unsupervised data augmentation method called syntax fine-tuning (Qi et al., 2023), knowledge preservation fine-tuning (Siddiq et al., 2023a), and task-oriented fine-tuning (Sun et al., 2023b), etc.
@@ 5.2. What prompt engineering techniques are applied to improve the performance of LLMs in SE tasks?
Prompt engineering is a method of enhancing model performance by using task-specific instructions, known as prompts, without modifying the core model parameters. This approach enables LLMs to seamlessly integrate into downstream tasks solely based on the given prompts, guiding model behavior without the need to update model parameters (Sahoo et al., 2024). Fig. 9 presents eight prompt engineering techniques currently applied in the LLM4SE domain.
Figure 9. The prompting engineering techniques used in LLMs for SE tasks. See Appendix C for the full table including references.
Few-shot prompting. Few-shot prompting involves providing a limited number of examples or instructions to the model to perform a specific task. The model learns from these examples and generalizes to similar tasks with minimal training data. In the surveyed LLM4SE research, 88 studies utilized few-shot prompting (Ahmed et al., 2024; Feng and Chen, 2023; First et al., 2023; Geng et al., 2024; Kang et al., 2022; Wei et al., 2022a; Xu et al., 2024; Zhang et al., 2023g). For instance, Geng et al. (Geng et al., 2024) adopted an in-context learning paradigm and providing a specific number of prompts simultaneously significantly outperformed state-of-the-art supervised learning methods in generating comments with multiple intents.
Zero-shot prompting. In zero-shot prompting (Radford et al., 2019), the model is expected to perform a task without any explicit training on that task. Instead, it relies on the prompt provided during inference to generate the desired output. Following few-shot prompting in terms of usage frequency, 79 studies adopted zero-shot prompting (Kou et al., 2023a; Li et al., 2023e; Luo et al., 2022; Pei et al., 2023; Schäfer et al., 2023b; Weyssow et al., 2023b; Wu et al., 2023a; Yan et al., 2023a). For example, Li et al. (Li et al., 2023e) introduced CodeEditor, a pre-trained model specifically designed for code editing, and demonstrated its effectiveness in automatic code editing under zero-shot settings.
Chain-of-Thought (CoT) prompting. Wei et al. (Wei et al., 2022b) introduced a prompting technique called Chain-of-Thought (CoT), which involves each prompt building upon the preceding one, resulting in a coherent chain of reasoning that enhances the model’s ability to generate well-structured and thoughtful responses. Huang et al. (Huang et al., 2023g) proposed a novel method leveraging the fault-tolerance and comprehension capabilities of pre-trained LLMs to generate Control Flow Graphs. This method involves a Chain-of-Thought (CoT) with four steps: structural hierarchy extraction, nested code block extraction, CFG generation for nested code blocks, and merging of CFGs for all nested code blocks. Tian et al. (Tian and Chen, 2023) also introduced the first test case-driven code generation technique, named TCoT, to further enhance LLMs’ capabilities in code generation. Including the two studies mentioned earlier, a total of 18 studies applied CoT to improve LLMs’ performance in SE tasks (Deng et al., 2023d; Feng and Chen, 2023; Huang et al., 2023g, f; Li et al., 2023m, d; Liu et al., 2024b; Mu et al., 2023).
Automatic Prompt Engineer (APE). Inspired by classical program synthesis and human prompt engineering methods, Zhou et al. (Zhou et al., 2023b) introduced an Automatic Prompt Engineer (APE) for automatic instruction generation and selection. APE is a system designed to automatically generate effective prompts for LLMs based on the desired task. It aims to simplify the process of prompt engineering by automating the generation of task-specific instructions. Sharing a similar concept of automated prompts, Sun et al. (Sun et al., 2023b) proposed a new prompt learning framework called PromptCS. PromptCS trains a prompt agent that can generate continuous prompts to fully explore LLMs’ potential in code summarization tasks. Continuous prompts, generated under the guidance of LLMs, are easier for LLMs to comprehend compared to manually written discrete prompts.
Chain of Code (CoC) prompting. CoC prompting (Li et al., 2023g) is similar to CoT prompting but is specifically tailored for programming tasks. It involves providing a sequence of prompts or code snippets to guide the model’s code generation process. Huang et al. (Huang et al., 2023a) proposed CodeCoT, and Le et al. (Le et al., 2023) proposed CodeChain, both of which are reasoning frameworks that better guide LLMs in code generation.
Automatic Chain-of-Thought (Auto-CoT) prompting. Auto-CoT (Zhang et al., 2022d) is an automated version of CoT prompting where the sequence of prompts is generated automatically based on the input and desired task. Paranjape et al. (Paranjape et al., 2023) introduced a framework, Automatic. ART, for generating intermediate reasoning steps automatically. ART can select multi-step reasoning and tools from a task library based on given tasks at any time and has been experimentally proven effective in code tasks.
Modular-of-Thought (MoT) prompting. In code generation tasks, LLMs often generate solutions in the form of a single block of code, limiting their effectiveness in handling complex problems. To overcome this limitation, Li et al. (Li et al., 2023a) proposed the Modular-of-Thought Coder (MoTCoder). They introduced a new MoT prompting optimization framework to facilitate task decomposition into logical subtasks and submodules. Experimental results demonstrate that MoTCoder significantly improves the modularity and correctness of solutions generated by LLMs in programming tasks.
Structured Chain-of-Thought (SCoT) prompting. Considering that source code contains rich structural information, Li et al. (Li et al., 2023d) proposed SCoT prompting specifically for code generation tasks. Researchers enable LLMs to use program structure to construct CoTs (i.e., intermediate natural language reasoning steps) to obtain SCoTs. Then, LLMs generate the final code based on SCoTs. Compared to CoT prompts, SCoT prompts explicitly constrain LLMs to consider how to address requirements from the source code perspective. Evaluations across multiple benchmarks show that SCoT significantly enhances LLMs’ performance in code generation.
In addition to the eight prompting techniques mentioned above, we identified 76 studies where researchers, although not explicitly mentioning the application of any of the aforementioned prompting techniques, carefully designed prompts or proposed new strategies based on prompts to apply LLMs to SE tasks better. For instance, Ren et al. (Ren et al., 2023) proposed a code generation method based on knowledge-driven prompt chains. Li et al. (Li et al., 2023m) applied differential prompting to ChaGPT to better identify test cases that cause failures in buggy programs. Ahmed et al. (Ahmed et al., 2024) enhanced the performance of LLMs in code summarization tasks using automatic semantic augmentation prompts.
@@ 5.3. How are evaluation metrics utilized to assess the performance of LLM4SE tasks?
Evaluating the performance of LLM4SE is a crucial aspect of their development and deployment (Kang et al., 2022). Benchmarking against existing datasets and using baselines are common practices to evaluate the effectiveness of LLMs (Cassano et al., 2023). However, given the diversity of SE tasks, a single evaluation metric may not suffice to capture the model’s performance comprehensively. Thus, researchers often employ a range of evaluation metrics tailored to specific problem types (Mastropaolo et al., 2021b; Niu et al., 2022; Salza et al., 2022). We categorize the SE tasks summarized from 395 papers into four categories according to their addressed problem types, i.e., regression, classification, recommendation, and generation tasks, as displayed in Fig. 10 (b). The selection of evaluation metrics depends on the target problem types. For example, MAE (Mean Absolute Error) has been used for regression tasks (Fu and Tantithamthavorn, 2022). We summarize the most frequently used evaluation metrics for each task type.
@@
Problem Type Metric Total
Regression MAE (Mean Absolute Error) (1) 1
Classification Precision (35) Recall (34) 147
F1-score (33) Accuracy (23)
AUC (Area Under the ROC Curve) (9) ROC (Receiver Operating Characteristic) (4)
FPR (False Positive Rate) (4) FNR (Falser Negative Rate) (3)
MCC (Matthews Correlation Coefficient) (2)
Recommendation MRR (Mean Reciprocal Rank) (15) Precision/Precision@k (6) 39
MAP/MAP@k (6) F-score/F-score@k (5)
Recall/Recall@k (4) Accuracy (3)
Generation BLEU/BLEU-4/BLEU-DC (62) Pass@k (54) 338
Accuracy/Accuracy@k (39) EM (Exact Match) (36)
CodeBLEU (29) ROUGE/ROUGE-L (22)
Precision (18) METEOR (16)
Recall (15) F1-score (15)
MRR (Mean Reciprocal Rank) (6) ES (Edit Similarity) (6)
ED (Edit Distance) (5) MAR (Mean Average Ranking) (4)
ChrF (3) CrystalBLEU (3)
CodeBERTScore (2) MFR (Mean First Ranking) (1)
PP (Perplexity) (1)
  • *See Appendix D for the full table including references.
@@
For classification tasks, the most commonly used metrics are Precision (Biswas et al., 2020; Chen et al., 2023a; Ezzini et al., 2022; Fatima et al., 2022; He et al., 2022), Recall (Biswas et al., 2020; Chen et al., 2023a; Ezzini et al., 2022; Fatima et al., 2022; He et al., 2022; Hey et al., 2020) and F1-score (Alhamed and Storer, 2022; Biswas et al., 2020; Chen et al., 2023a; Ezzini et al., 2022; Fatima et al., 2022; He et al., 2022), with 35, 34, and 33 tudies, respectively, employing these metrics. For example, in the study conducted by Khan et al. (Khan et al., 2021), F1-score is utilized to evaluate the performance of an automatic bug-fixing model. Similarly, Sharma et al. (Sharma et al., 2022) use Precision and Recall to assess the effectiveness of a transformer-based model for code summarization. These metrics are essential for evaluating the model’s ability to correctly classify code snippets (Fatima et al., 2022) or identify specific SE properties (Chen et al., 2023a).
@@
For recommendation tasks, MRR (Mean Reciprocal Rank) is the most frequent metric, used in 15 studies (Ciborowska and Damevski, 2022; Izadi et al., 2022; Li et al., 2021; Lin et al., 2021; Rahmani et al., 2023; Salza et al., 2022; Shi et al., 2022; Wei et al., 2022a). MRR is employed to measure the effectiveness of recommendation systems for code completion, as demonstrated in the study by Ciborowska et al. (Ciborowska and Damevski, 2022). Precision@k (He et al., 2023; Ciborowska and Damevski, 2022; Lin et al., 2021; Zhu et al., 2023) and F1-score@k (He et al., 2023; Lin et al., 2021; Zhu et al., 2022, 2023) are also utilized in recommendation tasks, with 6 studies each. These metrics are used to evaluate the precision and F1-score of the recommended code snippets or code completions.